OpenStudy (alina123):

your baby cousin was born at a height of 1 1/12 ft. over the next 5 years she grew to be a height of 3 1/2 ft. find the average yearly growth

OpenStudy (alina123):

Wouldn't you divide 3.5 by 17/12

jigglypuff314 (jigglypuff314):

I was thinking 3.5 minus 13/2 then divide that by 5 because subtracting like that would give you how much she grew over the 5 years then dividing it by 5 would give you the rate at which she grows (aka average yearly growth)

OpenStudy (alina123):

Where did you get 13 over 12 from

jigglypuff314 (jigglypuff314):

1 1/12 = 12/12 + 1/12 = 13/12

OpenStudy (alina123):

Oh yeh

OpenStudy (alina123):

The answer is 29 over 60

jigglypuff314 (jigglypuff314):

that's what I got :)

OpenStudy (alina123):

K I have two sheets that I'll post as a pdf and I need help with like 6 questions can you help me with that

jigglypuff314 (jigglypuff314):

I can try :) but I might have to leave in the middle?

OpenStudy (alina123):

Well wherever we get to

OpenStudy (alina123):

I'll inform you